OpenStudy (anonymous):

What is the value of -3 x (7 = (-3)

OpenStudy (anonymous):

the equation has a problem?

OpenStudy (solomonzelman):

Is it \[(-3) \times 7 = -3\] or\[-3x \times 7 = -3\]or \[-3x^7=-3\]can you write the question clearly please?

OpenStudy (solomonzelman):

You wrote\[-3 \times (7=(-3)\]

OpenStudy (anonymous):

SZ you left me :(

OpenStudy (anonymous):

What is the value of -3 x (7

OpenStudy (solomonzelman):

\[-3 \times 7?~~~~~then~~~-21\]

OpenStudy (anonymous):

What is the value of -3 x (7=(-3^2)divided by 3 sorry here is the right one